a senior class is planning a trip to Chicago. a bus seats 51 people and a van seats 8. if there are 150 students going om the trip in a total of 8 vehicles how many buses and vans are taken on the trip?
Answer by ewatrrr(24785)


Hi
total of 8 vehicles
Let x and (8-x) represent the number of buses and vans respectively
Question states***
51x + 8(8-x) = 150
Solving for x
43x + 64 = 150
43x = 86
x = 2, the number of buses. 6 vans (8-2)
CHECKING our Answer***
2*51 + 6*8 = 102 + 48 = 150